OpenStudy (anonymous):

Solve for x: x-7/3=10

OpenStudy (anonymous):

add 7/3 to both sides x= 12 1/3

OpenStudy (anonymous):

37

OpenStudy (anonymous):

Is your question supposed to be:\[\frac{ x - 7 }{ 3 } = 10\]? Because if that is the how your question is supposed to look, then you do have the right answer, 37 Good job!